Learning how to graph
2x+y=-5
3x+3y=9

Answer by mananth(16946)

2x+y=-5
3x+3y=9
Find the x & y intercept points by plugging x=0,y=0 in both equations.

equation I
when x= 0 y= -5 ,( 0 , -5 )
when y = 0 x= -5 ,( -2.5 , 0 )
Equation 2
when x= 0 y= 3 ( 0 , 3 )
when y = 0 x= 3 ( 3 , 0 )



You will observe that the point of intersection is -8 , 11
The solution of the system is ( -8 , 11 )